OpenStudy (anonymous):

are the ratios 8/12 and 1/25 proportional? explain

OpenStudy (michele_laino):

explanation: you have to calculate the minimum common multiple between (mcm) 12 and 25. since 12=3*2^2 and 25=5^2 then mcm=25*4*3=300

OpenStudy (michele_laino):

now 300/12=25 and 25*8=200, then your first fraction is equivalennt to: \[\frac{ 300 }{ 300 }\]

OpenStudy (michele_laino):

so we can write: \[\frac{ 8 }{ 12 }=\frac{ 200 }{ 300 }\]

OpenStudy (michele_laino):

since 300/25=12 and 12*1=12, so second fraction is equivalent to: \[\frac{ 12 }{ 300 }\]

OpenStudy (michele_laino):

namely we can write: \[\frac{ 1 }{ 25 }=\frac{ 12 }{ 300 }\]

OpenStudy (michele_laino):

finally since: \[\frac{ 200 }{ 300 }\neq \frac{ 12 }{ 300 }\] we conclude that: \[\frac{ 8 }{ 12 }\neq \frac{ 1 }{ 25 }\]

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

finally since 25 and 12 have no divisors in common, then our fractions can not be proportional each other

OpenStudy (michele_laino):

thanks!